OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (anonymous):

2-1 --- = 1/2 -4+6

OpenStudy (anonymous):

1. You first need the slope formula by solving, you will get = 1/2 2. Choose a point and use the point slope formula to get your equation y = 1/2*x + 4